OpenStudy (anonymous):

What is the equation of a line, in general form, with a slope of -2 and a y-intercept of 8? x + 2y - 8 = 0 2x + y - 8 = 0 2x - y + 8 = 0

OpenStudy (anonymous):

y=-2x+8

OpenStudy (anonymous):

y = mx + b m = slope b = y intercept -2x -y + 8 = 0

OpenStudy (anonymous):

-2x -y+8=0

OpenStudy (anonymous):

@xXstopXx and @Mesopretty are correct